# MISS: A Generative Pretraining and Finetuning Approach for Med-VQA

### v1

@@ 3.4 Tranfer and Caption
The purpose of TransCap is to overcome the current challenges in medical image-text datasets, which often contain a large amount of noise since their images and captions are mostly extracted from open-source papers. In the medical field, some unimodal tasks, such as medical image classification and lesion segmentation, often have high-quality open-source data. TransCap aims to utilize these unimodal datasets to construct high-quality multimodal image-text pairs.
@@
Fig. 2: Transfer and Caption unimodal images. We construct image descriptions based on image attributes and ChatGPT.
@@
TranCap defines six attributes for medical images: modality, plane, shape, size, organ, location, and pathology. As shown in Figure 2, for each input image $I$, TranScap constructs a corresponding dictionary $dict_{I}\{\}$ with keys representing the six image attributes. For unimodal medical image datasets, TranScap uses ChatGPT to generate attribute content based on dataset information and task labels. Then, it uses this attribute content as a prompt to input the LLM and requests it to generate multiple ways of expressing the attribute’s corresponding textual description. The attribute textual description serves as the value corresponding to the attribute key in $dict_{I}\{\}$. During pretraining, TranScap constructs captions by randomly sampling the various attribute contents of the input image dictionary. In this way, we can make use of the previously overlooked large amount of high-quality open-source unimodal image data and obtain captions that are more in line with human expression habits through LLMs. More details will be described in the appendix.
3.5 VQA Finetuning
The Med-VQA task requires generating answers to given questions and images. Previous Med-VQA tasks have been treated as classification or rank tasks, making the practical application value of models severely limited. To our knowledge, our model is the first fully generative model in the medical field (both training and testing stages are generative tasks).
@@ 3.5 VQA Finetuning
As shown in Figure 1(b), during the finetuning stage, the image input $I$ undergoes image encoding to extract image embeddings $\{i_{<cls>},i_{1},...,i_{n}\}$. The question input $Q$ is encoded by the JTM encoder to obtain question embeddings $\{q_{<encod>},q_{1},...,q_{n}\}$ and then interacted with the $\{i_{<cls>},i_{1},...,i_{n}\}$ in the cross-attention layer to obtain joint feature representations $f\in\mathbbm{R}^{n+1}$. The tokenized answer is then sent to the casual attention layer to obtain answer embeddings $\{a_{<decod>},a_{1},...,a_{n}\}$, which serve as query vectors in the cross-attention layer. These then interact with the joint feature representations, and the final output is used to calculate LM loss $\mathcal{L}_{LM}$ like Bert with the ground truth.
@@ 4 Experiment
In this section, we compare MISS with a series of previous medical VQA models. Since most methods treat VQA as a simple classification task or rank task, while our model treats VQA as a generation task, and most methods use different training paradigms and pretraining data scales, a direct result comparison may be unfair for our method. We also conduct extensive ablation studies on our method. Next, we will introduce our dataset, comparative experiments, and ablation studies. The implementation details and baseline of the experiments will be introduced in the appendix.
Table 1: Comparsion with other works which have different methods, pretraining images, training paradigm and types of task.
Methods Pretrain # images Training paradigm Type of task VQA-RAD SLALKE
@@
Basic Vision-Language Models (parameters <1B)
MEVF[16] 11,779 Meta Learning classification 75.1 43.9 - - - -
MMQ[15] - Supervised learning classification 75.8 53.7 67 - - -
VQAMIX[17] - Supervised learning classification 79.6 56.6 70.4 - - -
AMAM[23] - Supervised learning classification 80.3 63.8 73.3 - - -
PUBMEDCLIP-MEVF [19] - Pretraing-finetuning classification 78.1 48.6 66.5 76.2 79.9 77.6
CPRD[18] 22,995 Mean Teacher classification 80.4 61.1 72.7 83.4 81.2 82.1
MTL[24] 87,952 Pretraing-finetuning classification 79.8 69.8 75.8 86.1 80.2 82.5
M3AE[13] 298,000 Pretraing-finetuning classification 83.4 67.2 77 87.8 80.3 83.2
MUMC[12] 387,000 Pretraing-finetuning ranking 84.2 71.5 79.2 - - 84.9
ours (base model) 38,800 Pretraing-finetuning generating 80.35 71.81 76.05 82.91 81.47 82
Large Vision-Language Model (parameters >1B)
LLaVA (7B)[26] - Pretraing-finetuning ranking 65.07 50.00 - 63.22 78.18 -
LLaVA-Med (7B)[25] 1M Pretraing-finetuning ranking 84.19 61.52 - 85.34 83.08 -
LLaVA-Med (13B)[25] 1M Pretraing-finetuning ranking 81.98 64.39 - 83.17 84.71 -
4.1 Dataset
We consider two Med-VQA benchmarks: the VQA-RAD dataset[14] and the Slake dataset[21]. VQA-RAD contains 315 radiology images and 3,515 QA pairs annotated by clinicians, which are evenly distributed over the head, abdomen, and chest. SLAKE is a semantically-labeled knowledge-enhanced dataset for Med-VQA, which consists of 642 radiology images and 14,028 QA pairs created by experienced physicians. Details of both datasets will be introduced in the Appendix. VQA-RAD doesn’t provide a test set, and we extract 1,797 QA pairs to train the model, and the rest 451 pairs are used to test. For Slake, 14,028 QA pairs are divided into 70% training, 15% validation, and 15% testing subsets.
@@ 4.2 Comparison with Other Methods
We conduct a comparative evaluation of our method against the existing approaches on VQA-RAD and Slake. Compared to past research, our model is the only one that treats Med-VQA as a generative task, while others have approached VQA as answer classification or ranking tasks. Since generated evaluation metrics such as BELU cannot intuitively represent the model’s effectiveness in Med-VQA, we have relied solely on answer accuracy as our primary evaluation metric. For closed-ended questions with only “yes” or “no” answers, we utilize automated evaluation methods. For open-ended questions, we follow the approach outlined in [14] and conduct manual evaluations comparing the generated responses with ground-truth answers. We take the model without TransCap as our base model.
Fig. 3: Answers of our method and the ground truth.
Table 1 demonstrates our comparison with existing methods on Slake and VQA-RAD. Even if the current Large Vision-Language Model achieves better results, the extremely low pre-training cost and parameter amount of the base Vision-Language Model cannot be ignored. Apart from the large-scale vision-language model LLaVA (7B or 13B), our model has the smallest pre-trained image scale compared with methods in the pre-trained fine-tuning paradigm, using only 38,800 images. Nevertheless, our base model achieves the best accuracy in open-ended questions for Slake, which reaches 81.47%, surpassing all methods employing answer classification and ranking tasks. For VQA-RAD, although the test sets selected by each method may vary, the results show that our model still achieves good performance.
@@ 4.3 Abslation Studies
To demonstrate the effectiveness of different components of our method, we perform an ablation study on Slake, with the results shown in Table 2. There are several observations drawn from the results. The model without pretraining achieves only 50.99% accuracy on closed-ended questions, indicating that it can understand the task type through VQA fine-tuning, but cannot fully understand the semantics of medical images. When MISS did not utilize the JTM encoder and conventional multi-modal models were used to set up the text encoder and multi-modal encoder, our global accuracy rate was 1.64% lower than that of the base model, indicating that the JTM encoder can extract joint features more efficiently.
When our model uses both the JTM encoder and the TransCap method, we compare the impact of TransCap on our model by increasing the amount of pretraining data. As shown in the table, when using the TransCap method, with only an increase of less than 12k pretraining images, our open-ended accuracy and closed-ended accuracy increased by 1.03 and 0.5, respectively, demonstrating the positive effect of TransCap on VQA performance. Since most of the captions generated by TranCap are judgmental statements, the proportion of judgmental captions in the pretraining data continues to increase, resulting in a slight decrease in accuracy on open-ended questions; at the same time, it also leads to a certain increase in overall accuracy, ultimately reaching 83%.
@@
Methods Pretrain # images SLALKE
@@ 5 Conclusion
In this paper, we propose a pretraining and finetuning framework for medical VQA tasks. We treat medical VQA as a generative task and propose a Joint Text-Multimodal encoder and align image-text features through multi-task learning. Furthermore, we propose a Transfer-and-Caption method that extends the feature space of single-modal image datasets using LLMs, enabling those traditional medical vision-field task data to be applied to VLP. We demonstrate excellent results with fewer multimodal datasets and the advantages of generative VQA models through experiments. We hope that our method will encourage the development of Med-VQA in both data and model aspects.
Appendix A Appendix
A.1 Overview
In the appendix, we will supplement the content that cannot be described in detail due to space limitations in the main text. We will introduce several contents in turn: first, we will introduce the basic principle of Mask Language Modeling(MLM) loss and its calculation method in the pretraining and finetuning process of this article; then, we will use practical examples to specifically explain how the TransCap method proposed in this article constructs multi-modal image-text pairs based on LLM and unimodal image data; after that, we will supplement the implementation details and baselines of our experiments.
A.2 Details of MLM loss
In the MLM task of Bert, 15% of the tokens are randomly selected, 80% of them are replaced with a special token [MASK], 10% are randomly replaced with other words, and the other 10% are left unchanged. The model does not know whether the position is [MASK] or an original token or a random word, so it forces the model to predict the masked token according to the context. In our proposed paradigm, while pretraining the model, the tokenized caption embeddings $\{w_{<decod>},w_{1},...,w_{n}\}$ and the image embeddings $\{i_{<cls>},i_{1},...,i_{n}\}$ fuse in the cross-attention layer and while finetuning it’s joint feature $f$ who fuses with question embeddings $\{q_{<encod>},q_{1},...,q_{n}\}$. Both of them then are sent to the feed-forward layer and calculate MLM loss $\mathcal{L}_{MLM}$.
MLM will minimize a cross-entropy loss between predicted results and ground-truth results:
${\mathcal{L}_{MLM}=\mathbb{E}_{(I,\hat{T})\sim D}H(y^{msk},p^{msk}(I,\hat{T})}$ (4)
A.3 Details of TransCap
Transfer and Caption is a method based on large language models(LLMs) to extend the feature space of unimodal image data, which has never been explored in the medical field before. Specifically, medical images have a large amount of high-quality open-source unimodal data in other unimodal tasks, such as image classification and lesion segmentation. The labels of these data contain the attributes of medical images. We define the attributes of medical images as follows: modality, plane, shape, size, organ, location, and pathology. By obtaining the attribute content of images, we input it into LLMs and obtain different ways of describing all attribute content. By freely combining multiple attribute descriptions, we can obtain diverse text descriptions of a certain image. Below, we introduce the process of constructing image-text pairs using the TransCap method with a specific example.
The RSNA-PDC is a chest radiograph (CXR) dataset with a training set of 26,684 CXRs in three categories: Normal, No Lung Opacity/Not Normal, Lung Opacity. Based on the CXR attributes, the labels of each CXR image can be set to the following types: Modality (indicates data type), Class (can be: Normal, No Lung Opacity/Not Normal, Lung Opacity), Nums (indicates the number of lung opacities), Location (indicates the location of lung opacity). For example, one piece of data in the dataset is:
$dict${“Img_id”: “000db696-cf54-4385-b10b-6b16fbb3f985”, “Modality”: “CXR”, “Class”: “Lung Opacity”, “Nums”: 2, “Location”: “the upper left, the upper right”}
The attribute labels of each data will be input into LLM as a prompt, and it is required to generate a caption to describe the attributes of CXR. Additionally, to realize data enhancement, LLM will be utilized to rephrase, obtaining more captions with the same meaning, but more in line with human expression habits.
Ultimately, one-to-one caption sets for each CXR image are constructed, which contain multiple captions with different expressions. As mentioned in this paper, during pre-training, the captions in the caption set will be randomly sampled to construct an image-text pair, such as {“Img_id”: “000db696-cf54-4385-b10b-6b16fbb3f985”, “Caption”: “This chest X-ray reveals abnormalities in the lung, with the presence of 2 lung opacities located at the upper left, the upper right.”}
An image from an unimodal dataset is transformed into an image-text pair, and data enhancement is realized with the LLM to obtain a caption set which is more in line with human expression habits. Figure 4 compares the largest medical multimodal dataset, MedICat, collected from article centres, with image-text pairs generated by TransCap. It can be seen that the image-text pairs generated by the TransCap method contain less noise and are more humanized in terms of both image and caption. We believe that this method of combining LLMs to generate pretrained multimodal medical data will greatly encourage the development of this field.
A.4 Details of Experiments
Baseline: In this paper, we present an image-text pretraining framework tailored for generative tasks and propose a joint text-multimodal encoder to simultaneously extract features from both images and text through multi-task pretraining. To better compare the advantages of our approach with others, we have constructed a baseline for our study based on ALBEF. Specifically, in terms of model architecture, the baseline model comprises a vision transformer (ViT) base as the backbone for the image encoder, consisting of 12 transformer layers; a BERT-based text encoder and multimodal encoder, wherein the first six layers of the BERT model serve as the text encoder identical to the original BERT encoder, while the latter six layers incorporate cross-attention between the self-attention and feed-forward layers to function as the multimodal encoder; and a BERT-based text decoder, with each layer identical in structure to the BERT encoder.
In terms of the pretraining of the baseline model, it still follows ALBEF and sets up three pretraining tasks: Image-Text Contrastive Learning, Image-text Matching Learning, and Mask Language Modeling Learning. For fine-tuning the VQA task, we still use Mask Language Modeling Learning as the fine-tuning task. Since ALBEF still treats VQA as a RANK task in testing, we modified the output end of the model to enable it to directly output text. Finally, we constructed our baseline model based on the above methods, which is the method mentioned in the main text that does not utilize the JTM encoder.
Fig. 4: Comparison of data from MedICat Dataset and image-text pair data generated through TransCap. Image-text pairs generated by the TransCap method contain less noise and are more humanized in terms of both image and caption.
Pretraining Details: To better compare the superiority of our method and the impact of TransCap on our model, we choose to train a basic model that does not utilize the high-quality image-text pairs generated by TransCap but rather employs the MedICaT dataset used by most works in the literature. However, to demonstrate the model’s performance, we do not use the entire MedICaT dataset for training as in MAE and MUMC. Instead, we randomly select 38,800 radiological images from MedICaT to pretrain our model, a training scale that is only one-eighth to one-tenth of the training data used by other methods. At the same time, in the ablation experiments, to demonstrate the effectiveness of the TransCap method, we gradually incorporate the image-text pairs generated by TransCap into the basic model training, increasing the pretraining data size to 50,000, 70,000, and 90,000. Since TransCap mainly generates descriptive judgmental statements, Table 3 demonstrates that our TransCap method is effective in terms of the model’s performance on closed-ended questions.
Table 3: Ablation studies on different components of our method
Methods Pretrain # images SLALKE
CLOSED OPENED OVERALL
ours (base model) 38,800 82.91 81.47 82
baseline 38,800 82.82 79.11 80.36
ours (JTM+TranScap) 50,000 83.94 81.87 82.47
ours (JTM+TranScap) 70,000 83.94 81.44 82.38
ours (JTM+TranScap) 90,000 84.51 81.19 83
Implement Details: Here, we will present the experimental details of our pretraining and fine-tuning models. All of our training was conducted on a single NVIDIA RTX8000-50GB GPU. During the pretraining stage, we did not use any data augmentation techniques. We used the Adamw optimizer with cosine learning rate decay, an initial learning rate of 2e-5, weight decay of 0.05, a minimum learning rate of 0, and training for 100 epochs on the pretraining dataset. Additionally, during the pretraining stage, the input image size for our model was 224x224 pixels. In the fine-tuning stage, we used the same optimizer settings and learning rate as in the pretraining stage, with an input image size of 480x480 pixels, and trained for 200 epochs. For our baseline model, the Vit-based visual encoder consists of 12 layers of transformer, and both the JTM encoder and text decoder contain 12 layers of transformer-based layers.

### v3

@@ 3.4 Transfer and Caption
Transfer and Caption (TransCap) is a method based on Large Language Models (LLMs) to extend the feature space of unimodal image data, which has never been explored in the medical field before. The purpose of TransCap is to overcome the current challenges in medical image-text datasets, which often contain much noise since their images and captions are mostly extracted from open-source papers [25]. However, in the medical field, some unimodal tasks, such as medical image classification and lesion segmentation, often have high-quality open-source data. Our method aims to utilize these unimodal datasets to construct high-quality multimodal image-text pairs.
@@
Figure 2: Transfer and Caption unimodal images. We construct image descriptions based on image attributes and ChatGPT.
@@
TransCap defines six attributes for medical images: modality, plane, shape, size, organ, location, and pathology. As shown in Figure 2, for each input image $I$, TransCap constructs a corresponding dictionary $dict_{I}\{\}$ with keys representing the six image attributes. For unimodal medical image datasets, TransCap uses ChatGPT to generate attribute content based on dataset information and task labels. Then, it uses this attribute content as a prompt to input the LLM and requests it to generate multiple ways of expressing the attribute’s corresponding textual description. The attribute textual description serves as the value corresponding to the attribute key in $dict_{I}\{\}$. During pretraining, TransCap constructs captions by randomly sampling the various attribute contents of the input image dictionary. In this way, we can make use of the previously overlooked large amount of high-quality open-source unimodal image data and obtain captions that are more in line with human expression habits through LLMs.
For example, the RSNA-PDC [1] is a chest radiograph (CXR) dataset with a training set of 26,684 CXRs in three categories: Normal, No Lung Opacity/Not Normal, and Lung Opacity. Based on the CXR attributes, the labels of each CXR image can be set to the following types: Modality (indicates data type), Class (can be: Normal, No Lung Opacity/Not Normal, Lung Opacity), Nums (indicates the number of lung opacities), Location (indicates the location of lung opacity).
The attribute labels of each data will be input into LLM as a prompt, and it is required to generate a caption to describe the attributes of CXR. Figure 3 shows a series of examples where the attribute categories of the original single-modal dataset are translated into the caption of an image, the origin data on the left side shows the medical image attributes saved in dictionary form, and the right side shows the image caption constructed by the TransCap method.
Figure 3: Examples of constructing new image-text pair by TransCap. Discrete image attribute information is converted to image descriptions end-to-end by ChatGPT.
An image from an unimodal dataset is transformed into an image-text pair, and the multimodal data generation is realized with the LLM to obtain a caption set which is more in line with human expression habits. Figure 4 compares the largest medical multimodal dataset, MedICaT, collected from article centres, with image-text pairs generated by TransCap. One can observe that the image-text pairs gathered from open-source publications contain a lot of noise, such as blurry details and irrelevant captions, whereas those produced by the TransCap method are comparatively less noisy and appear more human-like in both image and caption.
3.5 VQA Fine-tuning
@@
As shown in Figure 1(b), during the fine-tuning stage, the image input $I$ undergoes image encoding to extract image embeddings $\{i_{<cls>},i_{1},...,i_{n}\}$. The question input $Q$ is encoded by the JTM encoder to obtain question embeddings $\{q_{<encod>},q_{1},...,q_{n}\}$ and then interacted with the $\{i_{<cls>},i_{1},...,i_{n}\}$ in the cross-attention layer to obtain joint feature representations $f\in\mathbbm{R}^{n+1}$. The tokenized answer is then sent to the causal attention layer to obtain answer embeddings $\{a_{<decod>},a_{1},...,a_{n}\}$, which serve as query vectors in the cross-attention layer. These then interact with the joint feature representations, and the final output is used to calculate LM loss $\mathcal{L}_{LM}$ like Bert with the ground truth.
Figure 4: Comparison of data from MedICat Dataset and image-text pair data generated through TransCap. Image-text pairs generated by the TransCap method contain less noise and are more humanized in terms of both image and caption.
@@ 4 Experiment
In this section, we compare MISS with a series of previous SOTA medical VQA models. Since most methods treat VQA as a simple classification task or rank task, while our model treats VQA as a generation task, and most methods use different training paradigms and pretraining data scales, a direct result comparison may be unfair for our method. We also conduct extensive ablation studies on our method to demonstrate the effectiveness of the JTM encoder and the TranCap. Next, we will introduce our baseline of the experiments, implementation details, dataset, comparative experiments, and ablation studies respectively.
4.1 Baseline
In this paper, we present an image-text pretraining framework tailored for generative tasks and propose a joint text-multimodal encoder to simultaneously extract features from both images and text through multi-task pretraining. To better compare the advantages of our approach with others, we have constructed a baseline model for our study, which is based on ALBEF[18]. Specifically, in terms of model architecture, the baseline model comprises a ViT-Base as the backbone for the image encoder, consisting of 12 transformer layers; a BERT-based text encoder and multimodal encoder, wherein the first six layers of it serve as the text encoder, which is identical to the original BERT encoder, while the latter six layers incorporate cross-attention between the self-attention and feed-forward layers to function as the multimodal encoder. A BERT-based text decoder is connected to the multimodal encoder and used for causal language modeling.
In terms of the pre-training of the baseline model, it still follows ALBEF and sets up three pre-training tasks: Image-Text Contrastive Learning, Image-text Matching Learning, and Mask Language Modeling Learning. For fine-tuning the VQA task, we still use Mask Language Modeling Learning as the fine-tuning task. Since ALBEF still treats VQA as a RANK task in testing, we modified the output end of the model to enable it to directly generate text.
Table 1: Comparsion with other works which have different methods, training paradigms and types of tasks on ACC (%). w/o” means the without.
Methods Training paradigm Type of task VQA-RAD SLALKE
@@ 4.1 Baseline
Small-scale Vision-Language Models
MEVF[23] Meta learning classification 75.1 43.9 - - - -
MMQ[6] Supervised learning classification 75.8 53.7 67 - - -
VQAMIX[10] Supervised learning classification 79.6 56.6 70.4 - - -
AMAM[24] Supervised learning classification 63.8 80.3 73.3 - - -
CPRD[21] Pretraing-finetuning classification 80.4 61.1 72.7 83.4 81.2 82.1
PUBMEDCLIP-MEVF[9] Pretraing-finetuning classification 78.1 48.6 66.5 76.2 79.9 77.6
M3AE[3] Pretraing-finetuning classification 83.4 67.2 77 87.8 80.3 83.2
MTL[4] Pretraing-finetuning classification 79.8 69.8 75.8 86.1 80.2 82.5
MUMC[20] Pretraing-finetuning ranking 84.2 71.5 79.2 - - 84.9
OURS(w/o Transcap) Pretraing-finetuning generating 80.35 71.81 76.05 82.91 81.47 82
Large-scale Vision-Language Model
LLaVA(7B)[15] Pretraing-finetuning generating 65.07 50.00 - 63.22 78.18 -
LLaVA-Med(7B)[16] Pretraing-finetuning generating 84.19 61.52 - 85.34 83.08 -
LLaVA-Med(13B)[16] Pretraing-finetuning generating 81.98 64.39 - 83.17 84.71 -
4.2 Dataset and Metrics
We consider two Med-VQA benchmarks: the VQA-RAD dataset[13] and the Slake dataset[22]. VQA-RAD contains 315 radiology images and 3,515 QA pairs annotated by clinicians, which are evenly distributed over the head, abdomen, and chest. SLAKE is a semantically-labeled knowledge-enhanced dataset for Med-VQA, which consists of 642 radiology images and 14,028 QA pairs created by experienced physicians. VQA-RAD doesn’t provide a test set, and we extract 1,797 QA pairs to train the model, and the rest 451 pairs are used to test. For Slake, 14,028 QA pairs are divided into 70% training, 15% validation, and 15% testing subsets.
Considering that existing state-of-the-art (SOTA) Med-VQA methods have different task paradigms, we cannot take BELU, BERTScore and et.al metrics to evaluate our model, which is usually adopted to evaluate generative models. To intuitively compare with the existing methods, we still choose accuracy (ACC (%)) as the only evaluation metric, although the evaluation of ACC is not necessarily fair for us compared with the methods using the classification paradigm or ranking paradigm.
@@ 4.3 Comparison with Other Methods
We conduct a comparative evaluation of our method against the existing SOTA approaches on VQA-RAD and Slake. For small-scale VLMs, our model is the only one that treats Med-VQA as a generative task compared with past research, while others have approached VQA as answer classification or ranking tasks. When we evaluate the ACC, for closed-ended questions with only “yes” or “no” answers, we utilize automated evaluation methods. For open-ended questions, we follow the approach outlined in [13] and conduct manual evaluations comparing the generated responses with ground-truth answers. We take the model without TransCap as our base model.
Table 1 demonstrates our comparison with existing methods on the Slake Dataset and the VQA-RAD Dataset. Even if the current Large-scale VLM achieves better results, the extremely low pre-training cost and parameter amount of the small-scale VLM cannot be ignored. Apart from the large-scale VLM LLaVA (7B or 13B), our base model achieves the best accuracy in open-ended questions for Slake, which reaches 81.47%, surpassing all methods employing answer classification and ranking tasks. For VQA-RAD, although the test sets selected by each method may vary, the results show that our model still achieves good performance. Table 2 demonstrates the comparison with those adopting the pre-training and fine-tuning paradigm on the scale of images for pre-training, our model has a smaller pre-trained image scale compared with methods in the pre-trained fine-tuning paradigm, using only 38,800 images. The competitive results demonstrate the efficiency of the JTM encoder.
@@
Figure 5: Answers of our method and the ground truth (Label).
@@ 4.4 Abslation Studies
Table 2: Comparison of SOTA methods adopting pre-training and fine-tuning paradigm but with different numbers of pre-trained images on open-ended ACC.
Methods Pre-train # images Type of task VQA-RAD SLALKE
CPRD 22,995 classification 61.1 81.2
PUBMEDCLIP-MEVF 11,779 classification 48.6 79.9
M3AE 298,000 classification 67.2 80.3
MTL 87,952 classification 69.8 80.2
MUMC 387,000 ranking 71.5 -
OURS(w/o Transcap) 38,800 generating 71.81 81.47
LLaVA(7B) - generating 50.00 78.18
LLaVA-Med(7B) 1M generating 61.52 83.08
LLaVA-Med(13B) 1M generating 64.39 84.71
To demonstrate the effectiveness of different components of our method, we perform an ablation study on Slake, with the results shown in Table 3. There are several observations drawn from the results. The model without pre-training achieves only 50.99% accuracy on closed-ended questions, indicating that it can understand the task type through VQA fine-tuning but cannot fully understand the semantics of medical images. When MISS did not utilize the JTM encoder and conventional multi-modal models were used to set up the text encoder and multi-modal encoder, our global accuracy rate was 1.64% lower than that of the base model, indicating that the JTM encoder can extract joint features more efficiently.
When our model uses both the JTM encoder and the TransCap method, we compare the impact of TransCap on our model by increasing the amount of pretraining data. As shown in the table, when using the TransCap method, with only an increase of less than 12k pretraining images, our open-ended accuracy and closed-ended accuracy increased by 1.03 and 0.5, respectively, demonstrating the positive effect of TransCap on VQA performance. Since most of the captions generated by TransCap are judgmental statements, the proportion of judgmental captions in the pretraining data continues to increase, resulting in a slight decrease in accuracy on open-ended questions; at the same time, it also leads to a certain increase in overall accuracy, ultimately reaching 83%.
@@
Methods Pre-train # images SLALKE
@@
4.5 Implement Details
Here, we will present the experimental details of our pre-training and fine-tuning models. All of our training was conducted on a single NVIDIA RTX8000-48GB GPU. During the pre-training stage, we did not use any data augmentation techniques. We used the Adamw optimizer with cosine learning rate decay, an initial learning rate of 2e-5, weight decay of 0.05, a minimum learning rate of 0, and training for 100 epochs on the pre-training dataset. Additionally, during the pretraining stage, the input image size for our model was 224x224 pixels. In the fine-tuning stage, we used the same optimizer settings and learning rate as in the pre-training stage, with an input image size of 480x480 pixels, and trained for 200 epochs. For our baseline model, the Vit-based visual encoder consists of 12 layers of transformer, and both the JTM encoder and text decoder contain 12 layers of transformer-based layers.
@@ 5 Conclusion
In this paper, we propose a pre-training and fine-tuning framework for Med-VQA tasks. We treat Med-VQA as a generative task and propose a Joint Text-Multimodal encoder and align image-text features through multi-task learning. Furthermore, we propose a Transfer-and-Caption method that extends the feature space of single-modal image datasets using LLMs, enabling the traditional medical vision-field task data to be applied to VLP. We demonstrate excellent results with fewer multimodal datasets and the advantages of generative VQA models through experiments. We hope that our method will encourage the development of Med-VQA in both data and model aspects.
6 Acknowledgement
This work is supported in part by the National Key R&D Program of China (2021ZD011 3503) and in part by the Shanghai Municipal Science and Technology Committee of Shanghai Outstanding Academic Leaders Plan (No. 21XD1430300).